OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\Large \int \frac{xe^x}{(1+x)^2}~dx\]

OpenStudy (freethinker):

show us what you did

OpenStudy (freethinker):

start from the format of IBP

OpenStudy (dumbcow):

integration by parts... u = e^x dv = x/(1+x)^2 du = e^x v = ln(1+x) + 1/(1+x)

OpenStudy (freethinker):

hmmm

OpenStudy (dumbcow):

looks like you will need to repeat IBP

OpenStudy (luigi0210):

I think what I did was probably not the best way.. I had these values instead: \(u=xe^x\) and \(\Large dv=\frac{dx}{(1+x)^2}\) \(du=e^x+xe^x dx\) and \(\Large v=-\frac{1}{(1+x)}\)

OpenStudy (freethinker):

you can try u = x e^x du = e^x (x+1) using product rule dv = 1/(1+x)^2 dx v = -1/(x+1) use uv - integral(vdu)

OpenStudy (kainui):

I was thinking about how to do this a clever way by seeing that 1/(1+x)^2 is the derivative of the infinite geometric sum of (-x)^n. But I'm still working on it.

OpenStudy (freethinker):

correction: u = x e^x du = e^x (x+1)dx dv = 1/(1+x)^2 dx v = -1/(x+1)

OpenStudy (freethinker):

integral e^x dx-(e^x x)/(x+1)

Miracrown (miracrown):

It may help with this integral to begin with a plain U substitution. With this choice for U, the integral becomes the following:  thinking of a better way to do this integral... hmm